# Zero-Shot Image Denoising for High-Resolution Electron Microscopy

### v1

@@ IV-D1 Low-Dose Te Crystal STEM Data
IV-D2 ZSM-5 & MOR Zeolites STEM Data
To further apply this method, two kinds of zeolites, ZSM-5 zeolite and MOR zeolite were imaged under very low-dose conditions and then denoised. It is worth mentioning that zeolites, as one kind of import nano-porous materials that were widely used in catalysis, separation, etc, are extremely sensitive to electron beams. Thus, low electron dose conditions were applied, resulting in low SNR of STEM images which hinder the visualization of atomic structure.
IV-E Effectiveness of Randomness in Sub-sampler
@@
TABLE II: Comparisons of fix and random location sub-sampling strategy in sub-sampler. Quantitative results (PSNR/SSIM) are evaluated on simulated TEM datasets with three different noise levels.
@@
IV-F Ablation Study
IV-F1 Influence of Sampling stride in Random Sub-sampler
@@
IV-F2 Influence of the Sampling Times in Inference
@@
IV-F3 Influence of Encoder Networks
To further explore the influence of the encoder network on the performance of Noise2SR, we compare the U-net based encoder with two well-known transformer-based backbones Uformer [52] and Restormer [53]. Additionally, to better demonstrate the superiority of Noise2SR, we also compare the performance of Noise2Self with the three encoders and a SOTA BSN, FBI-Net [17].
@@ IV-F3 Influence of Encoder Networks
TABLE III: Quantitative comparisons of Noise2Self and Noise2SR with three distinct encoder networks and FBI-Net on two noisy samples. The computational complexity of different networks is also compared. MACs are computed on patch size of $128\times 128$.
@@
IV-F4 Influence of Different Upsampling Strategies
@@
Table IV shows the quantitative results. The results show that for the test sample with a PSNR of 3 dB, the denoising performance of the three upsampling methods is comparable, with Noise2SR using pixel shuffling showing a slight advantage. For the test sample with a higher SNR of 10 dB, Noise2SR using transposed convolution achieves a slight improvement of 0.3 dB compared to the other two methods. It is worth noting that the bilinear interpolation upsampling method, which does not have any trainable parameters, achieves comparable performance with the added benefit of faster inference time. This implies that Noise2SR is effective in denoising even when simpler upsampling techniques, such as bilinear interpolation, are employed.
@@
Upsampling 3 dB 10 dB Time
TransConv. 27.36/0.9292 36.22/0.9894 5.84s
PixelShuff. 27.51/0.9349 35.89/0.9881 5.48s
BiInter. 27.31/0.9325 35.59/0.9876 4.43s
V Conclusion
In this work, we propose Noise2SR, zero-shot self-supervised learning (ZS-SSL) method for denoising HREM images. We introduce the Random Sub-sampler to generate paired noisy images with different resolutions, preserving structural information while altering high-frequency signal-dependent noise. This enables effective noise removal through supervised training using paired noisy images. By employing a super-resolved strategy, the network generates denoised results of the original image size from sub-sampled noisy inputs. The generation of paired noisy images serves as data augmentation to mitigate overfitting in single image denoising. Furthermore, we present an approximate MMSE estimation of the clean signal by ensembling multiple denoised results from randomly sub-sampled noisy images of a single input. This further enhances the denoising performance. The training strategy and MMSE estimation enable Noise2SR to achieve superior denoising performance using only a single ultra-low SNR HREM image. Experimental results on synthetic and real HREM data demonstrate that Noise2SR outperforms state-of-the-art ZS-SSL methods and achieves comparable performance to supervised methods. Despite achieving state-of-the-art performance in zero-shot self-supervised denoising for HREM images, Noise2SR still has certain limitations. One such limitation is the random sub-sampling would introduce high-frequency information loss in the input noisy image. Consequently, when denoising images rich in details, Noise2SR with MMSE estimation tends to produce relatively smooth results. Another limitation is that the zero-shot DL denoiser demands minutes of training time, preventing it from achieving real-time denoising performance. Besides the limitations, We believe that this work has the potential to improve the signal-to-noise ratio of images in materials imaging domains.

### v2

@@ IV-E Comparison of Time Required for Zero-shot Denoiser
IV-F Effectiveness of Randomness in Sub-sampler
@@
TABLE II: Comparisons of fix and random location sub-sampling strategy in sub-sampler. Quantitative results (PSNR/SSIM) are evaluated on simulated TEM datasets with three different noise levels. The best performance is highlighted in bold, while the second-best performance is underlined.
@@
IV-G Ablation Studies
IV-G1 Influence of Sampling stride in Random Sub-sampler
@@
IV-G2 Influence of the Sampling Times in Inference
@@
IV-G3 Influence of Encoder Networks
To further explore the influence of the encoder network on the performance of Noise2SR, we compare the U-net based encoder with two well-known transformer-based backbones Uformer [50] and Restormer [51]. To better demonstrate the superiority of Noise2SR, we also compare the performance of Noise2Self with the three encoders and a SOTA BSN, FBI-Net [17]. Additionally, we compute the multiply-accumulate operations (MACs) of different networks to compare their computational complexity.
@@ IV-G3 Influence of Encoder Networks
Furthermore, we compared training and inference time for Noise2SR with three encoder networks. The quantitative comparison is shown in Fig. 13. The results indicate that adopting a transformer-based encoder network increases training and inference time. Especially Regformer, it takes nearly five times longer than U-Net.
TABLE III: Quantitative comparisons of Noise2Self and Noise2SR with three distinct encoder networks and FBI-Net on two noisy samples. The computational complexity of different networks is also compared. Multiply-accumulate operations (MACs) are computed on a patch size of $128\times 128$. The best performance is highlighted in bold, while the second-best performance is underlined.
@@
Figure 13: Quantitative comparison of training and inference time of Noise2SR with three encoder networks. The inference time is computed on denoising an image with a size of $1024\times 1024$.
IV-G4 Influence of Different Upsampling Strategies
@@
Table IV shows the quantitative results. The results show that for the test sample with a PSNR of 3 dB, the denoising performance of the three upsampling methods is comparable, with Noise2SR using pixel shuffling showing a slight advantage. For the test sample with a higher SNR of 10 dB, Noise2SR using transposed convolution achieves a slight improvement of 0.3 dB compared to the other two methods. The results indicate that Noise2SR is relatively robust to various upsampling strategies, including bilinear interpolation, which has no training parameters. This flexibility allows for selecting upsampling strategies based on the desired trade-off between performance and speed in different application scenarios.
@@
Upsampling 3 dB 10 dB
TransConv. 27.36/0.9292 36.22/0.9894
PixelShuff. 27.51/0.9349 35.89/0.9881
BiInter. 27.31/0.9325 35.59/0.9876
IV-H Failure Cases Illustrations
We present two failure cases of Noise2SR for a comprehensive evaluation and a better understanding of its limitations. We test the performance of Noise2SR on denoising simulated TEM images that are extremely corrupted by severe Poisson-Gaussian noise with $(a=1,b=0.02)$ and $(a=2,b=0.02)$. To more directly assess the severity of image corruption, we calculated the SNR of the noisy image using the following formula:
$\text{SNR}_{\text{dB}}=10\log_{10}\left(\frac{\mathbb{E}[\mathbf{s}^{2}]}{\mathbb{E}[\mathbf{n}^{2}]}\right),$ (13)
where $\mathbf{s}$ and $\mathbf{n}$ denote the clean signal and noise respectively. Fig. 15 shows the failure cases. Although some edges of the atoms are still recoverable in the denoised images, where almost no information is visible in the original noisy images, the background of the denoised results exhibits significant artifacts. Additionally, the outlines of the atoms, particularly at the edges, become very unclear, as highlighted by the yellow arrows.
Figure 15: Illustration of failure cases: quantitative and qualitative results of Noise2SR performance on denoising simulated noisy TEM image corrupted by Poisson-Gaussian noise with $(a=1,b=0.02)$ and $(a=2,b=0.02)$.
We analyzed the potential reason for the failures could be: 1) The severe Poisson-Gaussian noise violates the self-supervised denoising assumption that noise is uncorrelated with the signal; 2) the small local patches in the noisy image may lack sufficient statistical information for effective denoising. To address these issues, future work could explore novel sampling or noise transformation strategies to reduce the correlation between signal and noise. Additionally, designing a network with a larger receptive field or incorporating multi-scale sub-sampled noisy images might improve performance.
V Conclusion & Discussion
In this work, we propose Noise2SR, zero-shot self-supervised learning (ZS-SSL) method for denoising HREM images. We introduce the Random Sub-sampler to generate paired noisy images with different resolutions, preserving structural information while altering high-frequency signal-dependent noise. This enables effective noise removal through supervised training using paired noisy images. By employing a super-resolved strategy, the network generates denoised results of the original image size from sub-sampled noisy inputs. The generation of paired noisy images serves as data augmentation to mitigate overfitting in single image denoising. Furthermore, we present an approximate MMSE estimation of the clean signal by ensembling multiple denoised results from randomly sub-sampled noisy images of a single input. This further enhances the denoising performance. The training strategy and MMSE estimation enable Noise2SR to achieve superior denoising performance using only a single ultra-low SNR HREM image. Experimental results on synthetic and real HREM data demonstrate that Noise2SR outperforms state-of-the-art ZS-SSL methods and achieves comparable performance to supervised methods. Despite achieving state-of-the-art performance in zero-shot self-supervised denoising for HREM images, Noise2SR has certain limitations. One such limitation is that random sub-sampling can lead to a loss of high-frequency information in noisy input images, causing Noise2SR to produce oversmooth results in images with rich detail. Additionally, Noise2SR may fail to produce satisfactory results under extreme signal-correlated noise. Another limitation is that the zero-shot DL denoiser demands minutes of training, preventing it from achieving real-time denoising performance. Despite these challenges, we believe that this work holds the potential to improve the signal-to-noise ratio in materials imaging domains.